# A Midsummer Meme’s Dream: Investigating Market Manipulations in the Meme Coin Ecosystem

### v1

@@ 4.2.2. Price Change Distribution
If we exclude inactive tokens (those with a stable price), Fig. 3 suggests that investing in meme coins could result in positive growth in about 50% of cases, sometimes yielding significant profits. However, this analysis overlooks two key factors. First, it does not account for the 361 tokens for which updated price data was unavailable—many of which likely lost all their value. Second, as discussed in Sec. 4.1.3, meme coins with high returns may be subject to market manipulations that artificially inflate their perceived value. To explore this further, we examine the 707 tokens that exhibited exceptionally high returns (¿100%) after three months.
5. Artificial Growth Strategies
@@
This section analyzes high-performing meme tokens: First, anomalous activity patterns; then, evidence of manipulations designed to distort the market. In particular, in order to analyze potential manipulations, we collected OHLCV (Open, High, Low, Close, Volume) data with hourly granularity from our initial data collection in mid-October 2024 through mid-January 2025 using the Gecko Terminal APIs (GeckoTerminal, 2025).
5.1. Market Anomalies in High-Return Tokens
To better understand the behavior of the 707 meme coins that exhibit returns exceeding 100%, we investigate potential anomalies in their market characteristics. To support this study, we collected additional data from Gecko Terminal and DexScreener, including ownership concentration metrics and security scores from monitoring services such as GoPlus (GoPlus, 2025), De.Fi (De.Fi, 2025), and Soul Scanner (Scanner, 2024).
5.1.1. Ownership Concentration Anomalies
@@
Ownership concentration occurs when a small number of wallets control most of the token supply. This concentration poses significant risks for investors, as major holders can manipulate prices through coordinated trading activities and other manipulative practices, effectively controlling the token’s value at the expense of smaller investors who often lack visibility into these ownership structures. We detect severe ownership anomalies across many high-return meme coins, with various techniques employed sometimes to concentrate token control while creating an illusion of distributed ownership.
@@
Concentrated Holdings. We identify 385 tokens (54.46%) with an alarming concentration of ownership among a small number of holders. The top 10 holders controlled an average of 77.85% of these tokens, with a median of 87.01%. Following standards used by established blockchain security platforms such as GoPlus (GoPlus, 2025) and CertiK (CertiK, 2025), we identify problematic concentration when the top holders control more than 30% of a token’s supply. This is the case for 359 tokens (50.78% of high-return tokens).
@@
Bundle Buy Concealment. A strategy developers can use to obscure their token holdings is leveraging bundle buys (e.g., JITO bundles (Explorer, 2025)), an approach that allows them to group multiple transactions together and execute them in a specific order. This enables the creation of bundles that first add liquidity (depositing tokens into a liquidity pool) and then immediately swap tokens, ensuring trades are executed right after liquidity is added. By distributing the swapped tokens across multiple wallets, they can mask true ownership and create the appearance of distributed holdings. We observe 133 tokens (18.81%) utilizing bundle buys, with an average of 15.70% of their supply involved. As for the concentrated holdings, we consider a bundle buy an anomaly when it exceeded 30% of a token’s supply, which occurred in 20 cases (2.83%).
@@
Fresh Address Distribution. Another concealment technique is the use of multiple newly created addresses. This tactic allows token creators or top holders to give the illusion of decentralized ownership while maintaining control for coordinated manipulation. We detect eight projects where a significant portion of tokens (more than 30%) are held in fresh addresses, wallets with no prior blockchain history. On average, these addresses control 40.25% of the token supply, with a median of 38%.
@@
Strategic Airdrop Distribution. Airdrops involve distributing tokens for free to multiple addresses and are commonly used for legitimate promotional purposes. However, they can also serve as a concentration tactic. By strategically distributing tokens to multiple controlled wallets, token creators can create the illusion of decentralized ownership while maintaining effective control. The average percentage of tokens distributed via airdrops is 63.43%, with a median of 79.36%. In particular, 24 tokens (3.40%) have more than 30% of their total supply allocated by airdrops.
5.1.2. Honeypots
@@
Unlike ownership concentration anomalies which focus on token distribution, honeypots represent a trading activity anomaly. These fraudulent tokens appear legitimate but are designed with malicious code that prevents investors from selling after purchase, effectively trapping their funds. This deceptive structure allows creators to collect investment capital while ensuring investors cannot exit their positions, regardless of how much the token’s value drops. We find 28 meme coins flagged as honeypots by Go Plus Token Security (GoPlus, 2025), representing a significant risk to unwary investors seeking high returns in the meme coin market.
5.1.3. Anomaly Prevalence
@@
Tab. 3 provides a summary of the statistics of the anomalies detected. Of the 707 high-performing tokens, 412 meme coins (58.27%) exhibit at least one anomaly indicator. Among these, 27 tokens show two anomalies: 26 combine honeypot behavior with ownership concentration, while one displays both airdrop activity and bundle buys. The most common issue is ownership concentration, where the top 10 holders own over 30% of the supply, occurring in 87.14% of the detected anomalies. Concerning the other indicators, although they are less frequent, they still help in flagging suspicious tokens that might otherwise escape detection.
Table 3. Anomaly statistics for high-performing meme coins.
@@
[…]
The table shows the anomaly statistics for meme tokens with high returns, including the number and percentage of tokens affected by each type of anomaly.
@@
The prevalence of anomalies in high-performing tokens raises serious doubt on the authenticity of their market success. To establish more concrete proof of artificial market activity, the following two subsections examine manipulation techniques that directly impact token valuation and visibility.
5.2. Wash Trading Analysis and Detection
@@
Wash trading is a deceptive practice in which the same entities buy and sell tokens back and forth to create an illusion of market activity. This manipulation inflates the trading volume without real ownership changes or market risk, creating a false perception of popularity and demand (Le Pennec et al., 2021). Meme coins are especially vulnerable to this strategy, as their success often depends on social sentiment and perceived popularity (Nani, 2022).
5.2.1. Initial Detection Methodology
@@
To detect potential wash trading, we identify suspicious trading patterns where volume increased dramatically (¿500% compared to the previous day) while the price remained nearly unchanged (absolute price change ¡5%). This pattern contrasts with natural market behavior, where significant volume surges typically drive notable price movements. We select the 500% volume threshold to focus only on extreme trading activity that deviates from normal market patterns, while the 5% price stability criterion helps distinguish these suspicious activities from legitimate high-volume trading that would typically cause significant price changes. Applying this method to the OHLCV data, we detect 3,264 potential wash trading instances involving 377 tokens.
5.2.2. Advanced Detection Techniques
@@
To strengthen our wash trading detection and provide more concrete evidence, we use the Dune Analytics API (Dune, 2025) to collect detailed transaction data. This enables us to identify the specific traders (makers) involved in suspicious transactions on the days flagged by our initial methodology. We apply three different techniques to detect wash trading activities.
@@
Zero-Risk Position. Wash trading, by definition, involves zero-risk market manipulation where traders buy and sell without taking actual market risk. To detect this strong evidence of wash trading, we look for makers who bought and sold almost identical amounts of the same token on the same day. To account for slippage and transaction fees, we allow for a small 2% difference between buy and sell volumes, consistent with methodologies used in previous works (La Morgia et al., 2023a). This approach detects 958 manipulations (29.35% of suspected cases), 271 involved tokens (71.88% of suspected tokens), and 1,777 unique makers.
@@
Circular Volume. A strong indicator of wash trading is when nearly all trading volume comes from the same set of addresses acting as both buyers and sellers. Such a high concentration of volume among circular traders strongly suggests coordinated manipulation rather than genuine activity. Thus, we analyze instances where at least 99% of daily trading volume is generated solely by makers who both bought and sold the token on the same day. This technique identifies 219 (6.71%) suspicious activities involving 128 (33.95%) tokens and 670 unique makers engaged in circular trading.
@@
Persistent Manipulator Network. The presence of persistent manipulator networks indicates organized and sustained efforts to manipulate the market. To uncover potential collusion networks, we examine whether the same makers participated in multiple wash trading events for the same token, focusing specifically on identifying persistent manipulators who appeared in every detected wash trading day for a given token. This technique detects 37 manipulations involving 14 tokens and 55 unique makers.
@@
Tab. 4 summarizes the results of the wash trading detection analyses. By combining the three detection approaches, we identify 1,014 events presenting strong evidence of wash trading activities (31.07% of suspected instances) and 282 tokens affected (74.8% of suspected tokens). The blockchain distribution of the manipulated tokens reveals that most are on Solana, with 181 tokens, followed by BSC with 52, Ethereum with 28, and Base with 21 tokens.
Table 4. Wash trading detection results by methodology
@@
Zero-Risk Position 958 (94.48%) 271 (96.10%)
Circular Volume 219 (21.60%) 128 (45.39%)
Persistent Manipulators 37 (3.65%) 14 (4.96%)
Total Detected Cases 1,014 282
The overlap of the detection techniques reveals that most wash trading activities (75.44%) are uniquely identified by zero-risk position, while only smaller proportions are exclusively detected through circular volume (4.04%) or persistent manipulator (1.38%). In particular, some activities exhibit multiple patterns: 17.46% are flagged by both zero-risk and circular volume techniques, while only 0.69% display all three wash trading indicators simultaneously. At the token level, the zero-risk position detects 94.48% of the meme coins suspected of wash trading, suggesting this approach forms the foundation of most wash trading operations.
5.2.3. Wash Trading Characteristics
After detecting instances of wash trading, we examine their frequency, volume characteristics, and temporal distribution to understand how manipulators create artificial market interest.
Recurrent Manipulations. Affected tokens experience an average of 3.59 wash trading operations, with a median of 3 and a maximum of 22, a result consistent across all blockchains. This pattern suggests recurring manipulation rather than isolated activities. Multiple wash trading operations on the same token likely serve to sustain the illusion of legitimacy and create a perception of growing interest over time.
Volume Generated. Wash trading operations create dramatic spikes in trading volume compared to the previous day, showing a median increase of 1,759% and an average increase distorted by extreme outliers reaching $10^{8}$%. During wash trading events, the average absolute trading volume is \$6,174, with a median of \$830, presenting platforms-specific differences. BSC shows the lowest median volume (\$366), followed by Base (\$516) and Solana (\$928), while Ethereum records the highest median volume (\$1673). This pattern likely reflects Ethereum’s higher transaction fees, which make small-volume wash trades economically impractical compared to other platforms with lower gas costs.
@@ 5.2.3. Wash Trading Characteristics
Figure 4. Temporal distribution of wash trading activity per token, relative to its creation date, across different blockchains.
[…]
The figure shows a scatterplot of wash trading activities occurring on each meme coin since its creation date, with different colors representing different blockchains.
@@
Fig. 4 illustrates the timeline of wash trading activity for meme coins. Most tokens experience their first manipulation within a median of 85 days, with some targeted earlier: 6.03% (17) within their first week and 24.47% (69) within their first month. Specifically, Solana tokens are targeted earlier, with a median of just 62 days after creation, compared to medians of at least 101 days on other blockchains. Moreover, these wash trading campaigns often continue for extended periods, averaging 87 days with a median of 64 days, indicating prolonged and sustained manipulation. Only 39.36% (111) of manipulated tokens experience activity for 30 days or less. For tokens subjected to multiple manipulations, the average time between consecutive operations is 33 days, with a median of 16 days. Specifically, 19.92% (202) of wash trading events occur within at most one week, and 49.41% (501) occur within at most one month. This pattern holds across blockchains, except for BSC, which has a slightly shorter median interval of 9 days.
@@
Few Wash Traders. Wash trading activities typically involve very few participants, with an average of 2.84 actors. In particular, 513 operations (50.59%) were performed by a single actor. This concentrated manipulation pattern suggests coordinated market manipulation carried out by a small group of entities or even individuals, a trend also observed in previous studies on wash trading (La Morgia et al., 2023a).
@@
These repeated wash trading operations performed by a few entities on the same tokens over extended periods suggest a strategic purpose: creating trading volume to maintain the appearance of legitimate market activity. Specifically, wash trading can serve two strategic purposes. First, it helps tokens reach the minimum trading volume required for listing on established cryptocurrency aggregators (CoinGecko, 2025d; CoinMarketCap, 2025b), ensuring visibility to potential investors. Second, it increases the ranking of a meme coin on platforms such as DexScreener (Dexscreener, 2025), which sorts tokens by trading volume. This increased visibility attracts unsuspecting investors, who perceive the artificial volume as evidence of genuine market interest.
5.3. Liquidity Pool-Based Price Inflation (LPI)
While wash trading distorts market perception through artificial volume, we identify another form of manipulation that directly targets token price rather than activity metrics. We define this technique as Liquidity Pool-Based Price Inflation (LPI), a manipulation that exploits the mechanics of liquidity pools on decentralized exchanges. LPI leverages the constant product formula ($x\times y=k$) used by liquidity pools to compute token prices. Indeed, in low-liquidity pools, even small trades can trigger disproportionate price movements. Manipulators exploit this mechanism by executing strategic purchases that dramatically inflate a token’s apparent market value with minimal investment, creating a deceptive impression of rapid appreciation that attracts unsuspecting investors.
To illustrate this manipulation, consider a simplified scenario: A liquidity pool contains 1 ETH and 1000 MEME tokens, making each MEME worth 0.001 ETH. When a manipulator purchases 500 MEME tokens for 0.5 ETH, the pool composition changes to 500 MEME and 1.5 ETH. Following the constant product formula, the token’s price now is 0.003 ETH, a 200% increase achieved through just a single transaction.
5.3.1. LPI Detection

### v2

@@ 4.2.2 Price Change Distribution
If we exclude inactive tokens (those with a stable price), Fig. 3 suggests that investing in meme coins could result in positive growth, sometimes yielding significant profits. However, this analysis overlooks two key factors. First, it does not account for the 361 tokens for which updated price data was unavailable, many of which likely lost all their value. Second, as discussed in Sec. 4.1.2, meme coins with high returns may be subject to market manipulations that artificially inflate their perceived value. To explore this further, we examine the 707 tokens that exhibited exceptionally high returns (>100%) after three months.
5 Artificial Growth Strategies
@@
This section analyzes high-performing meme tokens: First, anomalous activity patterns; then, evidence of manipulations designed to distort the market. In particular, in order to analyze potential manipulations, we collected OHLCV (Open, High, Low, Close, Volume) data with daily granularity from our initial data collection in mid-October 2024 through mid-January 2025 using the Gecko Terminal APIs [gecko_terminal_api].
5.1 Market Anomalies in High-Return Tokens
To better understand the behavior of the 707 high-performing tokens, we investigate anomalies in their market characteristics. To identify these irregularities, we leverage external security APIs (e.g., GoPlus [go_plus], De.Fi [defi_scanner], and Soul Scanner [soul_scanner]) which aggregate risk signals such as hidden ownership concentration. This investigation serves as a foundational risk assessment. While the detected anomalies are strongly correlated with manipulative behavior, they represent warning signals rather than definitive proof of manipulation. We establish concrete evidence of manipulation through the detailed analyses presented in Sec. 5.2 and Sec. 5.3.
5.1.1 Ownership Concentration Anomalies
@@
Ownership concentration occurs when a small number of wallets control most of the token supply. This concentration poses significant risks for investors, as major holders can manipulate prices through coordinated trading activities at the expense of smaller investors who often lack visibility into these ownership structures. We detect severe ownership anomalies across many high-return meme coins, with various techniques employed sometimes to concentrate token control while creating an illusion of distributed ownership.
@@
Concentrated Holdings. We identify 385 tokens (54.46%) with an alarming concentration of ownership among a small number of holders. The top 10 holders controlled an average of 77.85% of these tokens, with a median of 87.01%. Following standards used by established blockchain security platforms such as GoPlus [go_plus] and CertiK [certik_scan], we identify problematic concentration when the top holders control more than 30% of a token’s supply. This is the case for 359 tokens (50.78% of high-return tokens).
@@
Bundle Buy Concealment. A strategy developers can use to obscure their token holdings is leveraging bundle buys (e.g., JITO bundles [jito_bundle]), an approach that allows them to group multiple transactions together and execute them in a specific order. This enables the creation of bundles that first add liquidity (depositing tokens into a liquidity pool) and then immediately swap tokens, ensuring trades are executed right after liquidity is added. By distributing the swapped tokens across multiple wallets, they can mask true ownership and create the appearance of distributed holdings. We observe 133 tokens (18.81%) utilizing bundle buys, with an average of 15.70% of their supply involved. As for the concentrated holdings, we consider a bundle buy an anomaly when it exceeded 30% of a token’s supply, which occurred in 20 cases (2.83%).
@@
Fresh Address Distribution. Another concealment technique is the use of multiple newly created addresses. This tactic allows creators and top holders to give the illusion of decentralized ownership while maintaining control for coordinated manipulation. We detect 8 projects where a significant portion of tokens (more than 30%) are held in fresh addresses, wallets with no prior blockchain history. On average, these addresses control 40.25% of the token supply, with a median of 38%.
@@
Strategic Airdrop Distribution. Airdrops involve distributing tokens for free to multiple addresses and are commonly used for legitimate promotional purposes. However, they can also serve as a concentration tactic. By strategically distributing tokens to multiple controlled wallets, token creators can create the illusion of decentralized ownership while maintaining effective control. The average percentage of tokens distributed via airdrops is 63.43%, with 24 tokens (3.40%) having more than 30% of their supply allocated by airdrops.
5.1.2 Honeypots
@@
Unlike ownership concentration anomalies, which focus on token distribution, honeypots represent a trading activity anomaly. These fraudulent tokens appear legitimate but are designed with malicious code that prevents investors from selling after purchase, effectively trapping their funds. This deceptive structure allows creators to collect investment capital while ensuring investors cannot exit their positions, regardless of how much the token’s value drops. We find 28 meme coins flagged as honeypots by Go Plus Token Security [go_plus], representing a significant risk to unwary investors seeking high returns in the meme coin market.
5.1.3 Anomaly Prevalence
@@
Tab. 3 provides a summary of the statistics of the anomalies detected. Of the 707 high-performing tokens, 412 meme coins (58.27%) exhibit at least one anomaly indicator. The most common issue is ownership concentration, where the top 10 holders own over 30% of the supply, occurring in 87.14% of the detected anomalies. Concerning the other indicators, although they are less frequent, they still help in flagging suspicious tokens that might otherwise escape detection.
Table 3: Anomaly statistics for high-performing meme coins.
@@
The prevalence of anomalies in high-performing tokens raises serious doubts about their market success. To establish more concrete proof of artificial market activity, the following two subsections examine manipulation techniques that directly impact token valuation and visibility.
5.2 Wash Trading Analysis and Detection
@@
Wash trading is a deceptive practice in which the same entities buy and sell tokens to create an illusion of market activity. This manipulation inflates the trading volume without real ownership changes or market risk, creating a false perception of popularity and demand [le2021wash]. Meme coins are especially vulnerable to this strategy, as their success often depends on social sentiment and perceived popularity [nani2022doge].
5.2.1 Initial Detection Methodology
@@
To detect potential wash trading, we identify suspicious trading patterns where volume increased dramatically (>500% compared to the previous day) while the price remained nearly unchanged (absolute price change <5%). This pattern contrasts with natural market behavior, where significant volume surges typically drive notable price movements. These thresholds are heuristics implementing definitions provided in prior works [le2021wash, victor2021detecting]. The 500% volume threshold focuses on extreme trading activity that deviates from normal market patterns, while the 5% price stability criterion helps distinguish these suspicious activities from legitimate high-volume trading that would typically cause significant price changes. Applying this method to the OHLCV data, we detect 3,266 potential wash trading instances involving 379 tokens.
5.2.2 Advanced Detection Techniques
@@
To strengthen our wash trading detection and provide more concrete evidence, we use the Dune Analytics API [dune_api] to collect detailed transaction data. This enables us to identify the traders (makers) involved in suspicious transactions on the days flagged by our initial methodology. We apply three different techniques to detect wash trading activities.
@@
Zero-Risk Position. Wash trading, by definition, is a zero-risk manipulation where traders buy and sell without taking actual market risk. Following the methodology established by [la2023game], we look for makers who bought and sold almost identical amounts of the same token on the same day. To account for slippage and transaction fees, we allow for a small 2% difference between buy and sell volumes, consistent with methodologies used in [la2023game]. This approach detects 989 manipulations (30.28% of suspected cases), 276 involved tokens (72.82% of suspected tokens), and 2,687 unique makers.
@@
Circular Volume. A strong indicator of wash trading is when nearly all trading volume comes from the same set of addresses acting as both buyers and sellers. Thus, inspired by the cyclical pattern analysis in [victor2021detecting], we analyze instances where at least 99% of daily trading volume is generated solely by makers who both bought and sold the token on the same day. This technique identifies 223 (6.83%) suspicious activities involving 131 (34.56%) tokens and 1,148 unique makers engaged in circular trading.
@@
Coordinated Manipulator Network. To uncover potential collusion networks, we examine whether the same makers participated in multiple wash trading events for the same token, focusing specifically on identifying coordinated manipulators who appeared in every detected wash trading day for a given token. This technique detects 42 manipulations involving 17 tokens and 178 unique makers.
@@
Tab. 4 summarizes the wash trading detection analyses. By combining the three approaches, we identify 1,043 events presenting strong evidence of wash trading (31.94% of suspected instances) and 287 tokens affected (75.73% of suspected tokens). The blockchain distribution of the manipulated tokens reveals that most are on Solana, with 183 tokens, followed by BSC with 52, Ethereum with 29, and Base with 23.
Table 4: Wash trading detection results by methodology.
@@
Zero-Risk Position 989 (30.28%) 276 (72.82%)
Circular Volume 223 (6.83%) 131 (34.56%)
Coord. Manipulators 42 (1.29%) 17 (4.49%)
Total 1,043 287
The overlap of the techniques reveals that most wash trading activities (74.11%) are uniquely identified by zero-risk position, while only a smaller proportion is exclusively detected through circular volume (2.78%). At the token level, the zero-risk position detects 96.17% of the coins suspected of wash trading, suggesting this approach is the foundation of most wash trading schemes.
5.2.3 Wash Trading Characteristics
After detecting instances of wash trading, we analyze their patterns to understand how wash traders operate.
Recurrent Manipulations. Affected tokens experience an average of 3.63 wash trading operations, with a median of 3 and a maximum of 23, a result consistent across all blockchains. This pattern suggests recurring manipulation rather than isolated activities. Multiple wash trading operations on the same token likely serve to sustain the illusion of legitimacy and create a perception of growing interest over time.
Extreme Volume Increases. Wash trading operations create dramatic spikes in trading volume compared to the previous day, showing a median increase of 1,772% and an average increase distorted by extreme outliers reaching $10^{8}$%. During wash trading events, the average absolute trading volume is \$10,385, with a median of \$859, presenting platform-specific differences. BSC shows the lowest median volume (\$394), followed by Base (\$553) and Solana (\$964), while Ethereum records the highest median volume (\$1673). This pattern likely reflects Ethereum’s higher transaction fees, which make small-volume wash trades economically impractical compared to other platforms with lower gas costs.
@@ 5.2.3 Wash Trading Characteristics
Figure 4: Temporal distribution of wash trading activity per token, relative to its creation date, across different blockchains.
@@
Fig. 4 illustrates the timeline of wash trading activity for meme coins. Most tokens experience their first manipulation within a median of 85 days, with some targeted earlier: 6.27% (18) within their first week and 24.27% (71) within their first month. Specifically, Solana tokens are targeted earlier, with a median of just 61 days after creation, compared to medians of at least 101 days on other blockchains. Moreover, these wash trading campaigns often continue for extended periods, averaging 86 days, indicating prolonged and sustained manipulation. Only 40.07% (115) of manipulated tokens experience activity for one month or less. For tokens subjected to multiple manipulations, the average time between consecutive operations is 33 days, with a median of 16 days. Specifically, 19.94% (208) of wash trading events occur within at most one week, and 50.24% (524) occur within at most one month. This pattern holds across blockchains, except for BSC, which has a slightly shorter median interval of 9 days.
@@
Few Wash Traders. Wash trading activities typically involve very few participants, with an average of 3.92 actors and 514 operations (49.28%) conducted by a single actor. This manipulation pattern suggests coordinated market manipulation carried out by a small group of entities or even individuals, a trend also observed in previous studies on wash trading [la2023game].
@@
These repeated wash trading operations performed by a few entities on the same tokens over extended periods suggest a strategic purpose: creating trading volume to maintain the appearance of legitimate market activity. Specifically, wash trading can serve two strategic purposes. First, it helps tokens reach the minimum trading volume required for listing on established cryptocurrency aggregators [coingecko_listing, coinmarketcap_listing], ensuring visibility to potential investors. Second, it increases the ranking of a meme coin on platforms such as DexScreener [dexscreener], which sorts tokens by trading volume. This increased visibility attracts unsuspecting investors, who perceive the artificial volume as evidence of genuine market interest.
5.3 Liquidity Pool-Based Price Inflation (LPI)
In Sec. 4.1.2, we identified meme coins reporting suspiciously high market cap despite having minimal liquidity. Investigation of these cases reveals a previously undocumented manipulation strategy that directly inflates token prices through exploitation of automated market maker (AMM) mechanics. We term this technique Liquidity Pool-Based Price Inflation (LPI) and provide its first formal characterization.
Definition. We define LPI as a price manipulation technique that exploits the constant product formula ($x\times y=k$) used by AMM-based liquidity pools to generate disproportionate price increases with minimal capital investment. In this formula, $x$ represents the quantity of one token (e.g., ETH) and $y$ represents the quantity of another token (e.g., a meme coin). The constant $k$ must be maintained across all trades, meaning any purchase that removes tokens from the pool necessarily increases the price of the remaining tokens. Critically, in low-liquidity pools, even small purchases significantly alter the token ratio, triggering exponential price increases. Manipulators exploit this property to inflate the price with negligible capital, creating a misleading impression of rapid growth that attracts unsuspecting investors.
Illustrative Example. Consider a liquidity pool containing 0.01 ETH and 100 MEME tokens, where $k=0.01\times 100=1$. At this state, each MEME token is worth 0.0001 ETH. When a manipulator purchases 50 MEME tokens, they must add enough ETH to maintain the constant product. To remove 50 MEME (leaving 50 in the pool), the calculation becomes: $50\times y=1$, thus $y=0.02$ ETH must remain in the pool. Since the pool originally contained 0.01 ETH, the manipulator must add 0.01 ETH to acquire 50 MEME tokens. After this transaction, the pool contains 50 MEME and 0.02 ETH, establishing a new price of 0.0004 ETH per MEME, a 300% price increase achieved through a single transaction costing just 0.01 ETH (currently about \$30-40).
5.3.1 LPI Detection